OpenStudy (lgbasallote):

\(\LARGE \int \frac{\sqrt{x^2 - 4}}{x} dx\) Note: Trigonometric Substitution I think i got as far as \(\LARGE 2\int \cot ^2 \theta d\theta\) though i do not know if it is on the right track

OpenStudy (apoorvk):

Ahaa. Hmm. here, you're basically staring at your answer. \[cot^2\theta = cosec^2\theta - 1\] so your integral becomes " \[\int cosec^2\theta.d\theta - \int1.d\theta\] The derivative of -coxecx anyone?

OpenStudy (apoorvk):

ofcourse the '2' prefixed. and that will read '-cosec x', not "-coxecx"

OpenStudy (lgbasallote):

that's it? o.O well is cot^2 on the right track?

OpenStudy (apoorvk):

Oh that I don't know, I was assuming you did it right till there, let me check.

OpenStudy (anonymous):

put x=2sinu dx=2cosu du int (sqrt(x^2-4)/x)dx =int(sqrt(4sin^2 u -4) /2sinu) (2cosu)du =int(2cosu/2sinu)(2cosu)du =2int (cotu)(cosu)du : : :

OpenStudy (lgbasallote):

it came from a \(\LARGE \int \frac{tan \theta}{\sec\theta} \sec \theta \tan \theta d \theta\)

OpenStudy (apoorvk):

Niet. A wee wee bit of trouble. You did the substitution alright, but what did you do to 'dx' Did you make appropriate considerations for that? What substitution did you make?

OpenStudy (lgbasallote):

i used \(\LARGE x = 2\sec \theta\)

OpenStudy (lgbasallote):



OpenStudy (apoorvk):

Very well. so ...(am writing 'A' instead of 'theta' that's easier) dx = 2.secA.tanA.dA so, did you replace that instead of dx?

OpenStudy (lgbasallote):

yeah..as written in the above :P

OpenStudy (lgbasallote):

so...am i right with cot ^2 x?

OpenStudy (lgbasallote):

@apoorvk

OpenStudy (lgbasallote):

of course by x i meant theta

OpenStudy (apoorvk):

Oh okay you're doing it right. I'm drunk.

OpenStudy (lgbasallote):

ah i think i got it...

OpenStudy (lgbasallote):

im ready for your cutsies now =_=